15 rolls is what percent of 1000 rolls?
Answer by stanbon(75887)

15 rolls is what percent of 1000 rolls?
---------
Percent means "per 100".
---
x/100 = 15/1000
Cross-multiply to get:
x = 1500/1000
x = 1.5
Ans: 1.5%
===============
Cheers,
Stan H.